Kristin bought 1 pound of chocolate and 5 pounds of jelly beans for $9.24. Quwan bought 4 pounds of chocolate and 4 pounds of jelly beans for $18.40. How much does a pound of chocolate cost and how much does a pound of jelly beans cost?
Answer by stanbon(75887)

Kristin bought 1 pound of chocolate and 5 pounds of jelly beans for $9.24. Quwan bought 4 pounds of chocolate and 4 pounds of jelly beans for $18.40. How much does a pound of chocolate cost and how much does a pound of jelly beans cost?
------------
Equations:
c + 5j = 9.24
4c+ 4j = 18.40
-----------------------
Multiply thru the 1st equation by 4:
4c + 20j = 4*9.24
4c + 4j = 18.40
----
Subtract and solve for "J":
16j = 18.56
j = $1.16 (price of one lb of jelly beans)
---------------
Solve for "c":
c + 5j = 9.24
c + 5*1.16 = 9.24
c = 9.24-5.80
c = $3.44 (price of one lb of chodolate)
=========================
Cheers,
Stan H.
=========================